if dcs of two lines are (2/3,1/3,-2/3)&(3/7,-2/7,6/7)then the drs of the line bisecting the angle  between the line are

Grade:12

1 Answers

Nishant Vora IIT Patna
8 years ago
Dear student,

Let p=(2/3,1/3,-2/3) and q=(3/7,-2/7,6/7)

Now angle bisectors will be along
p cap + q cap
and
p cap – q cap

Thanks